# A Language-Agent Approach to Formal Theorem-Proving

### 11 Dec 2023 v2

@@ 2.2 A Markov Decision Process Formulation
By contrast, Copra is based on a view of automatic theorem-proving as a control problem. Like prior work on reinforcement learning (RL) for proof synthesis (Wu et al., 2021), we view a theorem-prover as a policy that interacts with a stateful proof environment (e.g., Lean) and model the interaction between the policy and the environment as a deterministic Markov Decision Process (MDP). We depart from prior RL-based work for theorem-proving by imposing a partial order on MDP states, adding execution feedback in error states, and allowing history-dependent policies.
Now we describe the different components of our proof MDP.
States. As before, let an obligation be a pair , where is a goal and a hypothesis. A state of the MDP is either a special symbol called or a set of obligations . The MDP has a unique initial state with a single obligation , where the goal and the hypothesis are extracted from the user-provided theorem that we are trying to prove. Its unique final state is the empty obligation set. The special symbol is accompanied by textual feedback in the form of an execution error message, execution feedback, from the proof environment.
@@
Actions and Transitions. The actions in our MDP are the proof environment’s tactics. The transition function determines the result of applying an action to a state . When is a tactic, we assume the underlying proof environment to return a state that results from applying to . If is a “bad” tactic, then equals ; otherwise, is a new set of obligations. We assume that our agent can evaluate for any state and action . While this assumption is unacceptable in many MDP problems, it is reasonable in the theorem-proving setting.
Rewards. As usual, we assume a reward function that evaluates an action at a state . Concretely, we consider rewards of the form , where is a very high positive value if , a high negative value if , and a small negative value otherwise. Intuitively, the small negative reward for successful action applications discourages the construction of overly long proofs when simpler proofs exist.
@@
A policy is a probabilistic function that maps histories to distributions over pairs , where is a state and is an action. Intuitively, at each point, the policy determines the next query to make to the proof environment. A policy can have an internal state as well as access to external knowledge (specifically, a lemma database). A trajectory of a policy is a history as above such that for each , . Letting each , the reward from a trajectory is simply the average . We define the aggregate reward of as the expected reward from trajectories sampled from .
Language Agents. Given our setup, one can naturally pose the problem of reinforcement-learning a policy with optimal aggregate reward. In this paper, we do not take on this problem. Instead, we consider a fixed policy — a wrapper around a pretrained LLM (GPT-4) that can learn in-context — and show that this policy can achieve a high reward. It is this policy that defines our language agent.
@@ 4 Evaluation
Implementing Copra. Our implementation of Copra can have GPT-3.5, GPT-4, GPT-4-turbo (OpenAI, 2023b) or CodeLlama (Roziere et al., 2023) as the underlying LLM and can interact with both the Lean and the Coq proof environments. Because of the substantial cost of GPT-4 queries, we cap the number of LLM queries that Copra can make at 60 for most of our experiments. At first, the retrieval mechanism is not used by the agent to keep the prompts smaller and cost-effective, but if the agent fails to find the proofs then retrieval is used to enrich the proof state before prompting the LLM. Similarly, informal proofs in the case of are only used when the agent fails to find the proof with and without retrieval. All the informal proofs are first generated using a separate one-shot prompt, and then Copra simply uses these proofs along with informal theorem statements to enhance its prompts. More details about the setup can be found in Section A.1.1.
The “system prompt” in the one-shot approach is slightly different than that for Copra, containing instructions to generate a proof in one go rather than step by step. For both Copra and the one-shot baselines, the prompt contains a single proof example that clarifies how proofs need to be formatted. This proof example remains the same for all test cases.
@@
Benchmarks. We evaluate our approach on two domains: (i) (Zheng et al., 2021), a collection of 244 Lean 3 formalizations of mathematics competition problems, solved using a range of techniques such as induction, algebraic manipulation, and contradiction; and (ii) a set of Coq problems from the CompCert compiler verification project (Leroy, 2009) that was previously used to evaluate the Proverbot9001 system Sanchez-Stern et al. (2020).
Baselines. We compare with one-shot invocations of GPT-3.5 and GPT-4 in both the and the Compcert domains. We also consider an ablation of Copra that uses GPT-3.5 as its LLM and another that does not use backtracking. For the dataset, we also have additional baselines with models like CodeLlama (Roziere et al., 2023), and ablations of Copra’s retrieval and use of informal proofs capabilities. As for fine-tuned baselines, a challenge is that all existing open-source theorem-proving systems only target a single-proof environment. As a result, we had to choose different baselines for the Lean () and Coq (Compcert) domains.
@@
Our fine-tuned baseline for the domain is ReProver, a state-of-the-art open-source prover that is part of the Leandojo project (Yang et al., 2023). We use BM25 search on Lean’s mathlib library for retrieval of relevant lemmas. We also use the informal theorem statements in the dataset first to generate all the informal proofs using the one-shot prompting of the underlying LLM. The prompt has one fixed example (for all invocations) of generating informal proof for a given informal theorem statement.
@@
In the Compcert domain, we compare with Proverbot9001 (Sanchez-Stern et al., 2020), which, while not LLM-based, is the best publicly available model for Coq. Unlike , this benchmark comes with a large training set as well as a test set, and we use the training set for retrieving relevant lemmas and definitions. Our retrieval mechanism, in this case, is a simple BM25 search.
Metric: pass@-guidance-steps. The standard metric for evaluating theorem-provers is pass@k (Lample et al., 2022; Yang et al., 2023). In this metric, a prover is given a budget of proof attempts; the method is considered successful if one of these attempts leads to success. However, a key objective of our research is to discover proofs quickly, with fewer LLM queries and lower wall-clock time. The pass@k metric does not evaluate this characteristic as it does not quantify the number of LLM queries or the amount of time needed by a proof attempt.
To address this concern, we introduce a new metric, pass@k-guidance, and evaluate Copra and its competitors using this metric. Here, we measure the number of correct proofs that a prover can generate with a budget of or fewer guidance steps from the LLM or any neural model. For LLMs, one guidance step is a single inference query. In our case, we use only a single attempt to prove the theorem, hence pass@k-guidance-step is comparable to pass@1 while also capturing finer details about the efficiency of the approach. One challenge here is that we want this metric to be correlated with the number of correct proofs that the prover produces within a wall-clock time budget; however, the cost of an inference query is proportional to the number of responses generated per query. To maintain the correlation between the number of inference queries and wall-clock time, we restrict each inference on LLM to a single response. (more details about the metric are in Section A.1.3).

### 13 Dec 2023 v3

@@ 2.2 A Markov Decision Process Formulation
By contrast, Copra is based on a view of automatic theorem-proving as a control problem. Specifically, we view a theorem-prover as a policy that interacts with a stateful proof environment (e.g., Lean) and model the interaction between the policy and the environment as a deterministic Markov Decision Process (MDP).
Proof MDP. Our proof MDP consists of a set of states, a set of actions, a deterministic transition function, a reward function, and an optional global context. The global context is a text string that describes the theorem to proven, as well as natural-language insights (Jiang et al., 2022a) about how to prove it. We describe the other components of the MDP below.
States. As before, let an obligation be a pair , where is a goal and a hypothesis. A state of the MDP is either a set of obligations , or an error state , where is a special symbol and is a textual error message. We denote the set of error states by .
The MDP has a unique initial state with a single obligation , where the goal and the hypothesis are extracted from the user-provided theorem that we are trying to prove. Its unique final state is the empty obligation set.
@@
Actions and Transitions. The actions in our MDP are the proof environment’s tactics. The transition function determines the result of applying a tactic to a state . If can be successfully applied at state , then is the new set of obligations resulting from the application. If is a “bad” tactic, then is an error state , where is the feedback from the proof environment that explains the error. We assume that we can evaluate for any state and action .
Rewards. As usual, we assume a reward function that evaluates an action at a state . Our rewards are of the form , where is a very high positive value if , a high negative value if , and a small negative value otherwise. Intuitively, the small negative reward for successful action applications discourages the construction of overly long proofs when simpler proofs exist.
@@
A policy is a probabilistic function that maps pairs , where is a history and is the global context, to distributions over pairs , where is a state and is a tactic. Intuitively, at each point, the policy decides the next tactic to apply.
A policy can have an internal state as well as access to external knowledge (e.g., lemma databases). A trajectory of a policy is a length- history as above such that for each , . Letting each , the reward from a trajectory is the average . The aggregate reward of is the expected reward from trajectories sampled from .
Language Agents. Given our setup, one can naturally pose the problem of reinforcement-learning a policy that obtains optimal aggregate reward, i.e., finds a short, correct proof to the theorem at hand. In this paper, we do not take on this problem. Instead, we consider a fixed policy — a wrapper around a pretrained LLM (e.g., GPT-4) that can learn in-context — and show that this policy can achieve a high reward. It is this policy that defines our language agent.
@@ 4 Evaluation
Benchmarks. We evaluate Copra on two domains: (i) (Zheng et al., 2021), a collection of 244 Lean 3 formalizations of mathematics competition problems, solved using a range of techniques such as induction, algebraic manipulation, and contradiction; and (ii) a set of Coq problems from the CompCert compiler verification project (Leroy, 2009) that was previously used to evaluate the Proverbot9001 system Sanchez-Stern et al. (2020).
Implementing Copra. Our implementation of Copra can have GPT-3.5, GPT-4, GPT-4-turbo (OpenAI, 2023b) or CodeLlama (Roziere et al., 2023) as the underlying LLM and can interact with both the Lean and the Coq proof environments. Because of the substantial cost of GPT-4 queries, we cap the number of LLM queries that Copra can make at 60 for most of our experiments.
@@
The Lean version of our system uses BM25 search on the mathlib library for retrieval of relevant lemmas. Inspired by the DSP approach (Jiang et al., 2022a), we also use the informal theorem statements in the dataset to generate informal proofs of theorems using one-shot prompting of the underlying LLM. The prompt has one fixed example (for all invocations) of generating informal proof for a given informal theorem statement. These informal proofs are included as part of the global context (we also consider an ablation where the informal proofs are not used).
@@
In the Compcert domain, we compare with Proverbot9001 (Sanchez-Stern et al., 2020), which, while not LLM-based, is the best publicly available model for Coq. Unlike , this benchmark comes with a large training set as well as a test set, and we use the training set for retrieving relevant lemmas and definitions. Our retrieval mechanism, in this case, is a simple BM25 search. Given that this domain does not come with informal theorem statements, we do not use informal proofs here.
At first, the retrieval mechanism is not used by the agent to keep the prompts smaller and cost-effective, but if the agent fails to find the proofs then retrieval is used to enrich the proof state before prompting the LLM. Similarly, informal proofs in the case of are only used when the agent fails to find the proof with and without retrieval. More details about the setup can be found in Section A.1.1.
Baselines. We compare with one-shot invocations of GPT-3.5 and GPT-4 in both the and the Compcert domains. The “system prompt” in the one-shot approach is slightly different than that for Copra, containing instructions to generate a proof in one go rather than step by step. For both Copra and the one-shot baselines, the prompt contains a single proof example that clarifies how proofs need to be formatted. This proof example remains the same for all test cases.
@@
We also consider an ablation of Copra that uses GPT-3.5 as its LLM and another that does not use backtracking. For the dataset, we also have additional baselines with models like CodeLlama (Roziere et al., 2023), and ablations of Copra’s retrieval and use of informal proofs capabilities. As for fine-tuned baselines, a challenge is that all existing open-source theorem-proving systems only target a single proof environment. As a result, we had to choose different baselines for the Lean () and Coq (Compcert) domains.
Our fine-tuned baseline for the domain is ReProver, a state-of-the-art open-source prover that is part of the Leandojo project (Yang et al., 2023). In the Compcert domain, we compare with Proverbot9001 (Sanchez-Stern et al., 2020), which, while not LLM-based, is the best publicly available model for Coq.
@@
Metric. The standard metric for evaluating theorem-provers is pass@k (Lample et al., 2022; Yang et al., 2023). In this metric, a prover is given a budget of proof attempts; the method is considered successful if one of these attempts leads to success. However, a key objective of our research is to discover proofs quickly, with fewer LLM queries and lower wall-clock time. The pass@k metric does not evaluate this characteristic as it does not quantify the number of LLM queries or the amount of time needed by a proof attempt.
To address this concern, we introduce a new metric, pass@k-guidance-steps, and evaluate Copra and its competitors using this metric. Here, we measure the number of correct proofs that a prover can generate with a budget of or fewer guidance steps from the LLM or any neural model. For LLMs, one guidance step is a single inference query. In our case, we use only a single attempt to prove the theorem, hence pass@k-guidance-step is comparable to pass@1 while also capturing finer details about the efficiency of the approach. One challenge here is that we want this metric to be correlated with the number of correct proofs that the prover produces within a wall-clock time budget; however, the cost of an inference query is proportional to the number of responses generated per query. To maintain the correlation between the number of inference queries and wall-clock time, we restrict each inference on LLM to a single response (see Section A.1.3 for more details).